OpenStudy (anonymous):

y-5/3 = 1 A) –2 B) 8 C) 18 D) 6 Can you tell me how to do it for future problems thanks!

OpenStudy (jiteshmeghwal9):

add \(\Large{\frac{5}{3}}\) both sides :)

OpenStudy (anonymous):

B) 8

OpenStudy (jiteshmeghwal9):

In future problems if u see any question similar to this then put all constants at one sides & all variables at one side by reversing their signs :)

OpenStudy (anonymous):

if you added the fraction to both sides wouldn't y = a fraction ;o? or decimal

OpenStudy (jiteshmeghwal9):

2x+3=5 Put all variables and their coefficients at one side & constants at one side by reversing their signs 2x=5-3 2x=2 x=1

OpenStudy (xishem):

Is this: \[\frac{y-5}{3}=1\]Or\[y-\frac{5}{3}=1\] ?

OpenStudy (arhin):

B) 8

OpenStudy (anonymous):

its the first one @Xishem

OpenStudy (jiteshmeghwal9):

If there is a plus sign then it will change to (minus) & if it is (minus) then it will change to (plus). If there is a multiplication sign then it will change to division & if division it will change to multiplication.

OpenStudy (anonymous):

would the answer still be one as the first problem in @xishems comment

OpenStudy (xishem):

\[\frac{y-5}{3}=1\]\[(3) \cdot \frac{y-5}{3}=1 \cdot (3)\]\[(5)+y-5=3+(5)\]\[y=8\]In an equation, if you perform the same operation to both sides, the equation stays true.

OpenStudy (xishem):

Notice how in the second line, the 3s on the left side cancel out and become y-5. And the +5 and -5 cancel out in the third line to make y. In solving for variables, you'll always want to perform operations such that this "cancellations" occur.

OpenStudy (anonymous):

o ty

OpenStudy (anonymous):

if two numbers times or divide each other they cancel out?